# LLMs Capture Urban Science but Oversimplify Complexity

### v2

@@ Abstract
Large language models (LLMs) are increasingly used in scientific reasoning and data generation tasks, yet their alignment with empirical knowledge and practical urban data in urban science remains unclear. Here, this study introduces AI4US (Artificial Intelligence for Urban Science), a framework for systematically evaluating five leading LLMs (ChatGPT, DeepSeek, Claude, Gemini, and ChatGLM) by generating urban data and testing their fidelity against real data distributions and the empirical results of three foundational theories, including urban scaling laws, urban distance decay, and urban vitality. LLMs reproduce core theoretical patterns with notable accuracy (average $\mathrm{R}^{2}=0.804$ for scaling; $0.988$ for decay), and exhibit capacity to simulate less formalized constructs such as urban vitality. However, generated data exhibit crucial limitations: poor diversity compared to real data; systematic deviations in fitted theoretical parameters (e.g., underestimated scaling exponents and decay rates); and limited improvement from prompt engineering. These findings suggest that while LLMs offer a promising research paradigm for exploring urban science, they fall short of capturing real-world complexity, underscoring the need for improved alignment and evaluation.
@@ 1 Introduction
Urban science has long sought to decipher the complex behaviors of cities from early demographic surveys to computational modeling [1, 2, 3, 4]. In recent decades, an explosion of urban data from remote sensing, sensor networks, and digital platforms has promised to revolutionize this endeavor [5, 6]. Yet, a fundamental paradox remains. Despite this data abundance, theoretical innovation has stagnated [7, 8, 9]. The field struggles to move beyond fragmented, context-dependent insights toward broadly generalizable urban theory. This impasse stems from a persistent methodological gap. The field lacks scalable and systematic experimental designs that allow for robust, cross-city replication and validation of scientific claims [8, 10, 11].
@@
This theoretical stagnation is rooted in the inherent difficulty of conducting controlled experiments in urban settings. Historically, urban experimentation has evolved through several paradigms, each with significant limitations. Early approaches treated the city as a living laboratory, relying on direct observation and policy trials, but these methods lacked scalability and control [12, 13]. Physical and analogue models offered controllable surrogates but were often oversimplified, failing to capture the non-linear dynamics of human behavior at a meaningful scale [14, 15]. The subsequent rise of digital models, such as cellular automata and agent-based simulations, introduced rule-based experimentation but hinged on hard-coded assumptions and static calibration, limiting their ability to represent the nuanced, adaptive behaviors of individuals [16, 17, 18]. These shortcomings are magnified by the profound transformations of contemporary cities, from autonomous mobility to platform economies, creating an urgent need for a new experimental paradigm capable of capturing the complex human-in-the-loop dynamics that define modern urban life [5, 19].
@@
Nowadays, advances in generative AI are spearheading a new paradigm of data-driven scientific discovery [20, 21, 22]. In fields like biology and materials science, its success lies in predicting specific physical states from vast solution spaces, exemplified by AlphaFold’s protein structure prediction [23] and MatterGen’s discovery of novel materials [24]. The challenge in urban science is fundamentally different. It is not to predict a discrete physical outcome [25], but to capture the broad statistical patterns that emerge from complex, adaptive urban systems, which are governed by context-dependent social behaviors rather than universal physical laws [26]. This distinction positions LLMs as a uniquely compelling tool. Trained on diverse records of human behavior and expression, LLMs offer a powerful pathway to generate synthetic urban data with high fidelity, enabling scalable urban experimentation where formal modeling has reached its limits.
@@
In urban science, however, the potential of generative AI to function as a new form of virtual city laboratory that can produce experimental data [27, 28], analyze urban phenomena, and drive theoretical advancement remains largely unexplored [20, 29]. Initial studies have shown promise, using LLMs to generate synthetic mobility data, simulate residents’ daily behaviors, or assess urban systems [22, 30]. However, these pioneering efforts often remain fragmented. A significant research gap persists in systematically validating the fidelity of these worlds generated by AI against established urban theories [31]. Furthermore, the inherent geographical and social biases within LLMs pose a critical challenge to their reliability as scientific instruments [32, 33]. Therefore, a structured and rigorous framework is necessary to move beyond isolated applications and systematically evaluate the potential and pitfalls of LLMs as a foundation for the next generation of urban science.
To address these gaps, we introduce the AI4US (Artificial Intelligence for Urban Science) framework. World models are internal, generative representations that simulate real-world dynamics [33]. This framework pioneers a methodology to evaluate LLMs as a first-generation proxy for urban world models. Our approach probes the implicit urban knowledge embedded within LLMs by testing their ability to reproduce three foundational urban theories: urban scaling laws, distance decay relationships, and Jane Jacobs’ urban vitality theory [29, 34, 35]. These theories were chosen to create a comprehensive testbed, as they represent fundamental urban mechanisms across different scales and dimensions: urban scaling laws govern macro level regularities, distance decay explains meso scale spatial interactions, and urban vitality captures micro scale human environment dynamics [36, 37, 38, 39, 40, 41]. Our findings reveal that while LLMs can replicate these theoretical patterns with notable fidelity, they also exhibit systematic limitations, underscoring that the path toward true theoretical discovery will require a new generation of world models capable of deeper causal reasoning.
@@ 2 Results
2.1 LLMs capture the fundamental forms of urban theories
Figure 1: Urban scaling law reproduced on generated data with LLMs: generated versus empirical coefficients. “Replicate Tests” contains 100 independent duplicate data generation experiments for each LLM, and each experiment generates data of 100 cities, among which GPT-4o has the best performance (see Supplementary Experiment 6). The diagram shows the results of GPT-4o.
Figure 2: Urban decay reproduced on generated data with LLMs: generated versus empirical statistical and spatial patterns. a, Concentric-ring land-density profiles for three representative cities on different continents across three decades (1990–2010), together with the corresponding “Mirage Cities” produced by GPT-4o when both city name and time period are included in the prompt. b, Sensitivity of the generated decay curves to prompts that specify different maximum radii and historical periods. c, Direct side-by-side comparison of Beijing’s empirical profile with that of its Mirage City generated by LLMs.
Figure 3: Urban vitality revealed on generated data with LLMs. “Replicate Tests” contains 100 independent duplicate data generation experiments for each LLM, and each experiment generates data of 100 blocks. The diagram shows the results of GPT-4o.
Figure 4: Data distribution divergence analysis. (a) Equal-width binning with relative frequency normalization was applied to convert raw sample counts into scale-invariant frequency distributions. Mean Absolute Error (MAE) of bin-level frequencies quantified the central tendency divergence between real data (R) and generated data (G), calculated as the sum of absolute differences across all 15 bins. Overlap Ratio (OR) measured quantile interval congruence between corresponding bins of R and G distributions. (b) Jensen-Shannon Divergence (JSD) comparisons among four datasets: real data (R), baseline generated data (G), China-prompted generated data (GC), and USA-prompted generated data (GA). The inclusion of geographical constraints, such as “China”, increased GPT-4o’s constraint-triggering rate by 24% (24/100 trials), e.g., “I can’t provide real-time data such as the current population”.
Figure 5: AI4US, the proposed framework utilizing LLMs to advance urban science.
@@ 2.1 LLMs capture the fundamental forms of urban theories
The investigation reveals that LLMs possess a remarkable ability to recognize and reproduce the foundational forms of established urban theories across multiple scales, from macro-level city systems to micro-level neighborhood dynamics. At the macro-scale, when tasked with generating data for urban scaling laws, LLMs consistently produce outputs that adhere to the characteristic power-law relationship. The generated data exhibits a strong fit to the scaling law equation, achieving a high average coefficient of determination ($R^{2}=0.804$) (Fig. 1a). Most models successfully replicated the expected super-linear relationship ($\beta>1$) for socioeconomic output and, in the case of GPT-4o, the sub-linear relationship for infrastructure. Moving to the meso-scale, the models display a non-trivial grasp of the concentric-zone structure posited by distance-decay theory. When generating land density surfaces, the outputs fit the inverse-S formulation with exceptional accuracy, showing a mean $R^{2}$ of 0.988 (Fig. 2a).
Furthermore, this capability extends beyond quantitative laws to less-formalized, qualitative theories. At the micro-scale, when confronted with urban vitality theory, LLMs demonstrate a clear alignment with its core tenets. By generating data for key indicators and then assessing their correlation with a livability score, the models produced relationships directionally consistent with Jane Jacobs’ theory (Fig. 3a). Specifically, indicators such as ’Building Mix Index’ and ’Tall Building’ showed the highest consistency with the theory, while ’Building Mix Index’ and ’Population Density’ demonstrated the most significant positive impact on the generated livability scores. In summary, LLMs can consistently generate data that is structurally ”form-similar” to a diverse set of urban theories, whether they are mathematically explicit or qualitative, establishing a baseline of theoretical competence.
@@ 2.2 Mirage cities systematically oversimplify real-world complexity
While LLMs can replicate theoretical forms, a deeper analysis reveals a significant gap between their generated ”Mirage cities” and the complexity of real urban systems. This discrepancy manifests as a series of systematic, non-random deviations. A primary finding is the lack of diversity in the generated data compared to empirical data. The analysis of data distributions reveals that LLM-generated data is significantly more homogeneous. Without specific geographic prompts, the generated data’s numerical scale diverges substantially from real data, resulting in a low average Overlap Ratio (OR) of just 28.6% per bin (Fig. 4a). More critically, the Jensen-Shannon Divergence (JSD) between real-world data samples (R-R) is, on average, 122.54% higher than the divergence within generated samples (G-G) (Fig. 4b). This indicates that real data exhibits substantially greater internal variation. This inherent ”over-smoothing” tendency is a foundational reason why LLM-generated data, while theoretically clean, lacks the real-world noise and heterogeneity essential for robust scientific analysis.
This lack of diversity contributes to specific parametric and structural distortions when fitting theories. At the macro-scale, a systematic underestimation of the scaling exponent ($\beta$) is observed in the urban scaling law test (Fig. 1b). The models consistently produce lower exponents for both socioeconomic output and infrastructure compared to established empirical values (e.g., $\beta\approx 1.15$ for GDP and $\beta\approx 0.85$ for infrastructure), suggesting they capture the ”shape” of the law but fail to grasp its true ”steepness” or magnitude. At the meso-scale, this manifests as the generation of ”idealized” spatial decay curves that are smoother than reality (Fig. 2a, Fig. 3b). The generated data gravitates toward the theoretical extrema, with values approaching 1 at the center and 0 on the periphery. Consequently, this idealized output fails to capture the ”messiness” of real cities, which contain features like central parks or suburban industrial belts that cause deviations from a perfect curve. The distributional difference in these generated decay curves is 149% smaller than that of empirical data, further underscoring their idealized nature.
@@ 2.3 The potential for improving AI-generated data
Given the observed deviations, this study investigated whether these gaps could be bridged through different forms of intervention. The results indicate that the fidelity of generated data possesses potential for improvement through both user interaction and model selection. Strategic prompting offers a viable pathway to enhance data fidelity. Prompting with specific geographic contexts demonstrates a clear effect on data distribution; when prompted with ”China,” the generated data showed a greater alignment with real data from Chinese cities, increasing the Overlap Ratio by 38.56% and decreasing the Mean Absolute Error by 52.14% (Fig. 4a, 4b). Similarly, at the meso-scale, prompting with parameters like city size or historical period directly influences the resulting decay curves (Fig. 2c). These results confirm that the model’s output is not static but sensitive to contextual cues, highlighting a clear potential for iterative refinement.
Furthermore, performance varies noticeably across the different LLMs evaluated. In the urban scaling law test, GPT-4o demonstrated the closest alignment with empirical results, while Claude 3.5 was most accurate in reproducing Zipf’s law. For distance decay, GPT-4o again showed more nuanced spatial reasoning, whereas models like ChatGLM-4 and DeepSeek-V3 tended to produce less diverse outputs. In the urban vitality assessment, GPT-4o consistently outperformed all other models. These variations suggest that factors intrinsic to the models, such as architecture, training data, and fine-tuning strategies, significantly influence the characteristics and fidelity of the generated urban data.
2.4 A new paradigm for qualitative theory exploration
The test of urban vitality, a theory that is descriptive and lacks a single mathematical form, reveals a distinct capability of LLMs, shifting their role from a ”physical simulator” to a ”socio-cultural simulator.” By prompting an LLM to act as a ”resident” evaluating a neighborhood, the task is not to solve a physical equation, but to have the model tap into its vast training data of human text and discourse to simulate a ”social consensus” on what makes a place livable. The model’s outputs revealed nuanced and theoretically consistent insights (Fig. 3). For instance, it identified ’Building Mix Index’ and ’Tall Building’ as having the highest consistency with vitality theory, while ’Aged Building’ showed the greatest deviation. In terms of impact, it judged ’Building Mix Index’ and ’Population Density’ as having the most significant positive effect on livability scores (Fig. 3b). This demonstrates that LLMs can serve as a powerful ”hypothesis generation engine.” They can process the complex, multi-faceted information embedded in qualitative theories and produce structured, testable hypotheses about the relative importance of different urban elements. This capability to ”operationalize” qualitative knowledge marks a significant new potential for AI in advancing urban science.

### v3

# GenAI Models Capture Urban Science but Oversimplify Complexity

@@ Abstract
Generative artificial intelligence (GenAI) models are increasingly used for scientific data generation, yet their alignment with empirical knowledge in urban science remains unclear. Here, we introduce AI4US (Artificial Intelligence for Urban Science), a framework that systematically evaluates leading GenAI models by testing their fidelity in generating both symbolic and perceptual urban data. For the symbolic domain, we benchmark generated data against foundational urban theories concerning scale, space, and morphology. For the perceptual domain, we validate the models’ visual judgments against human benchmarks and, critically, leverage their generative control to conduct in causal experiments on urban perception. Our findings show that while GenAI models reproduce core theoretical patterns, the generated data exhibit crucial limitations: poor diversity, systematic parametric deviations, and improvement from prompt engineering. To address this, we introduce a post-hoc calibration procedure using optimal transport, which produces synthetic symbolic datasets with demonstrably higher fidelity.
@@ 1 Introduction
Urban science has long sought to decipher the complex behaviors of cities from early demographic surveys to computational modeling [1, 2, 3, 4], which manifest across multiple scales (macro, meso, micro) and dimensions [5, 6]. The endeavor to formulate urban theory from empirical observation has traditionally relied upon two principal data modalities: symbolic data, which abstracts urban phenomena into text and numerics, and perceptual data, which captures the direct sensory experience of the urban environment through modalities like visual imagery [7, 8, 9]. Since the recent explosion of urban data [5, 6] from remote sensing, sensor networks, and digital platforms promised a revolution, struggling to move beyond fragmented, context-dependent insights toward broadly generalizable urban theory.
@@
However, theoretical innovation has stagnated [7, 8, 9]. This impasse stems from a persistent methodological gap: the lack of scalable and systematic experimental designs that allow for robust, cross-city replication and validation of scientific claims [8, 10, 11]. This theoretical stagnation is rooted in the inherent difficulty of conducting controlled experiments in urban settings. Historically, urban experimentation has evolved through several paradigms, each with significant limitations. Early approaches treated the city as a living laboratory, relying on direct observation and policy trials, but these methods lacked scalability and control [12, 13]. Physical and analogue models offered controllable surrogates but were often oversimplified, failing to capture the non-linear dynamics of human behavior at a meaningful scale [14, 15]. The subsequent rise of digital models, such as cellular automata and agent-based simulations, introduced rule-based experimentation but hinged on hard-coded assumptions and static calibration, limiting their ability to represent the nuanced, adaptive behaviors of individuals [16, 17, 18]. These shortcomings are magnified by the profound transformations of contemporary cities, from autonomous mobility to platform economies, creating an urgent need for a new experimental paradigm capable of capturing the complex human-in-the-loop dynamics that define modern urban life [5, 19].
@@
Nowadays, advances in GenAI are spearheading a new paradigm of data-driven scientific discovery [20, 21, 22]. In fields like biology and materials science, its success lies in predicting specific physical states from vast solution spaces, exemplified by AlphaFold’s protein structure prediction [23] and MatterGen’s discovery of novel materials [24]. The challenge in urban science is fundamentally different. AI’s role in urban science is to capture the broad statistical patterns emerging from complex, adaptive urban systems governed by context-dependent social behaviors rather than universal physical laws [25]. This distinction makes GenAI models particularly suitable. Their evolution from Large Language Models (LLMs), capable of generating and reasoning over symbolic data, to Multimodal Large Language Models (MLLMs) that also encompass perceptual data, unifies the two critical domains of urban information. Trained on diverse records of human behavior and expression, these models can thus generate high-fidelity synthetic data, enabling scalable urban experimentation where formal modeling has reached its limits.
@@
In urban science, however, the potential of GenAI to function as a new form of virtual city laboratory that can produce experimental data [26, 27], analyze urban phenomena, and drive theoretical advancement remains largely unexplored [20, 28]. Initial studies have shown promise, using GenAI models to generate synthetic mobility data, simulate residents’ daily behaviors, or assess urban systems [22, 29]. However, these pioneering efforts often remain fragmented. A significant research gap persists in systematically validating the fidelity of these worlds generated by GenAI against established urban theories [30]. Furthermore, the inherent geographical and social biases within GenAI models pose a critical challenge to their reliability as scientific instruments [31, 32]. Therefore, a structured and rigorous framework is necessary to move beyond isolated applications and systematically evaluate the potential and pitfalls of GenAI as a foundation for the next generation of urban science.
To address these gaps, we introduce the AI4US (Artificial Intelligence for Urban Science) framework. This framework establishes a ‘Generate-Evaluate-Calibrate’ workflow to systematically evaluate generative AI as a first-generation proxy by probing the implicit urban knowledge embedded within them. Our diagnostic approach evaluates the fidelity of AI-generated data across both symbolic and perceptual domains. For symbolic data, we test the models’ ability to reproduce foundational urban theories like urban scaling laws, distance decay, and urban vitality [28, 29, 30]. For perceptual data, we assess their multimodal capacity to generate realistic urban scenes and replicate human-centric visual perception. Crucially, this generative control enables a form of in causal experimentation, allowing for systematic tests of how specific urban elements impact human perception. These theories were chosen to create a comprehensive testbed, as they represent fundamental urban mechanisms across different scales and dimensions [31, 32, 33, 34, 35, 36]. Moving beyond diagnosis, this study introduces a novel post-hoc calibration procedure, demonstrating how a small sample of real data can be leveraged to rapidly assess these biases for a given urban dataset and subsequently correct the distribution of the larger generated dataset. This computationally efficient approach enhances the fidelity for downstream tasks, such as training predictive models or seeding simulations, without requiring costly modifications to the foundational GenAI models. Our findings reveal that the path toward the future will require advanced models [32].
@@ 2 Results
2.1 GenAI models capture the fundamental statistical forms of urban theories
@@
The investigation reveals that GenAI models possess a strong ability to recognize and reproduce the foundational forms of established urban theories across multiple scales, from macro-level city systems to micro-level neighborhood dynamics. At the macro-scale, when tasked with generating data for urban scaling laws, GenAI models consistently produce outputs that adhere to the characteristic power-law relationship. The generated data exhibits a strong fit to the scaling law equation, achieving a high average coefficient of determination ($R^{2}=0.804$) (Fig. 1a). Most models successfully replicated the expected super-linear relationship ($\beta>1$) for socioeconomic output and, in the case of GPT-4o, the sub-linear relationship for infrastructure.
Figure 1: Urban scaling law reproduced on generated data with GenAI models: generated versus empirical coefficients. ’Replicate Tests’ contains 100 independent duplicate data generation experiments for each GenAI model, and each experiment generates data of 100 cities, among which GPT-4o has best performance (See Supplementary Experiment 6). The diagram shows the results of GPT-4o.
Moving to the meso-scale, the models display a non-trivial grasp of the concentric-zone structure posited by distance-decay theory. When generating land density surfaces, the outputs fit the inverse-S formulation with high accuracy, showing a mean $R^{2}$ of 0.988 (Fig. 2a).
Figure 2: Urban decay reproduced on generated data with GenAI models: generated versus empirical statistical and spatial patterns. a, Concentric-ring land-density profiles for three representative cities on different continents across three decades (1990-2010), together with the corresponding ’Mirage cities’ produced by GPT-4o when both city name and time period are included in the prompt. b, Sensitivity of the generated decay curves to prompts that specify different maximum radii and historical periods. c, Direct side-by-side comparison of Beijing’s empirical profile with that of its mirage city generated by GenAI models.
Furthermore, this capability extends to descriptive theories that lack a single mathematical form, such as Jane Jacobs’ theory of urban vitality. The test of this theory reveals a distinct capability of GenAI, where the model is prompted to act as a ”resident” to simulate a ”social consensus” on what makes a place livable. The model’s outputs revealed nuanced and theoretically consistent insights, identifying ’Building Mix Index’ and ’Tall Building’ as having the highest alignment with the theory, while ’Aged Building’ showed the greatest deviation (Fig. 3). This demonstrates that an GenAI model can process the complex, multi-faceted information embedded in qualitative urban theories to generate data that is structurally ”form-similar” and directionally consistent with their core tenets. More importantly, it shows a capacity to translate these qualitative concepts into structured, testable hypotheses regarding the relative importance of different urban elements, establishing a baseline of theoretical competence.
Figure 3: Urban vitality revealed on generated data with GenAI models. ’Replicate Tests’ contains 100 independent duplicate data generation experiments for the GenAI models, and each experiment generates data of 100 blocks. The diagram shows the results of GPT-4o.
2.2 A new paradigm for qualitative theory exploration in visual urban space
This research demonstrates that GenAI models can function as a new paradigm for exploring and validating theories in visual urban space. The fidelity of the generated images was first confirmed in a blind expert validation test. AI-generated images received a mean realism score of 7.83 (SD = 1.6), which was not statistically different from the 7.68 (SD = 1.8) scored by real photographs (t(198) = 0.65, p = 0.52). This high fidelity was complemented by substantial diversity, achieved via a structured-prompting strategy that increased the average CLIP semantic distance to 0.3310. This represents a 35.8% improvement over a baseline single-prompt approach ($D_{CLIP}=0.2438$) and surpasses the diversity of a 1,000-image real-world benchmark ($D_{CLIP}=0.2905$), enabling the creation of semantically rich datasets.
Having established the quality of the generated data, we validated the AI’s ability to emulate human perception and identify its visual drivers. The AI’s perceptual judgments showed substantial agreement with human choices from the Place Pulse dataset across all six evaluated categories, with Kappa values consistently above 0.2 and agreement rates approaching 70% (Fig. 4a). Using this validated AI perception, we then quantified the influence of visual elements on these attributes through a multiple linear regression model. The analysis revealed several significant relationships (Fig. 4b). For instance, the presence of a ’person’ was a strong positive predictor for ’lively’ ($\beta=0.223,p<0.001$), while features such as ’wall’ negatively impacted the perception of ’beautiful’ ($\beta=-0.247,p<0.001$). To further validate these patterns, we compared our results to the benchmark findings of Place Pulse [37], which shows that ’tree’ and ’grass’ are strong positive predictors for ’beautiful’, while ’sky’ is a significant negative predictor for ’safe’.
Finally, the ”generate-modify-evaluate” paradigm was used to test the AI’s utility for causal inference by quantifying the impact of thematic interventions (Fig. 4c). The addition of Natural Elements produced the most pronounced effect, significantly increasing the perceived ’beautiful’ score by an average of 2.5 points ($\Delta$ Score) and having the strongest negative impact on the ’depressing’ score. Conversely, adding Traffic Elements was the most effective intervention for increasing the ’lively’ score, demonstrating the framework’s ability to test the perceptual outcomes of specific design strategies.
Figure 4: Potential of GenAI models in qualitative theory exploration in visual urban space. a, AI perception aligns with human judgment. A confusion matrix of pairwise choices (left) and per-category agreement rates with Cohen’s Kappa scores (right). b, Identifying influential visual elements. Standardized beta coefficients ($\beta$) from regression models reveal the key visual features that are positively or negatively correlated with each of the six perceptual scores. c, Causal inference via thematic interventions. The average change in perception scores ($\Delta$ Score) quantifies the causal impact of programmatically adding thematic element categories to scenes: Natural Elements (trees, grass, sky, water, bushes), Traffic Elements (cars, trucks, bridges, roads), and Built Elements (buildings, fences, walls). Error bars represent 95% confidence intervals.
@@ 2.3 Mirage cities systematically oversimplify real-world complexity
While GenAI models can replicate theoretical forms based on statistical data, a deeper analysis reveals a significant gap between their generated ”Mirage cities” and the complexity of real urban systems. This discrepancy manifests as a series of systematic, non-random deviations. A primary finding is the lack of diversity in the generated data compared to empirical data. The analysis of data distributions reveals that generated data is significantly more homogeneous. Without specific geographic prompts, the generated data’s numerical scale diverges substantially from real data, resulting in a low average Overlap Ratio (OR) of just 28.6% per bin (Fig. 5a). More critically, the Jensen-Shannon Divergence (JSD) between real-world data samples (R-R) is, on average, 122.54% higher than the divergence within generated samples (G-G) (Fig. 5b). This indicates that real data exhibits substantially greater internal variation. This inherent ”over-smoothing” tendency is a foundational reason why generated data, while theoretically clean, lacks the real-world noise and heterogeneity essential for robust scientific analysis.
This lack of diversity contributes to specific parametric and structural distortions when fitting theories. At the macro-scale, a systematic underestimation of the scaling exponent ($\beta$) is observed in the urban scaling law test (Fig. 1b). The models consistently produce lower exponents for both socioeconomic output and infrastructure compared to established empirical values (e.g., $\beta\approx 1.15$ for GDP and $\beta\approx 0.85$ for infrastructure), suggesting they capture the ”shape” of the law but fail to grasp its true ”steepness” or magnitude. At the meso-scale, this manifests as the generation of ”idealized” spatial decay curves that are smoother than reality (Fig. 2a, Fig. 3b). The generated data gravitates toward the theoretical extrema, with values approaching 1 at the center and 0 on the periphery. Consequently, this idealized output fails to capture the ”messiness” of real cities, which contain features like central parks or suburban industrial belts that cause deviations from a perfect curve. The distributional difference in these generated decay curves is 149% smaller than that of empirical data, further underscoring their idealized nature. This pattern of oversimplification extends to the visual domain. The AI’s perceptual judgment, while directionally correct, showed weaknesses in abstract cases, achieving lower consistency on subjective concepts like ’boring’ ($\kappa=0.155$) and ’lively’ ($\kappa=0.157$). Most notably, the AI exhibited a strong bias against ambiguity, agreeing with human ’equal’ judgments in only 2.1% of cases and instead forcing a definitive choice (Fig. 4a).
Figure 5: Data distribution divergence analysis. a, Equal-width binning with relative frequency normalization was applied to convert raw sample counts into scale-invariant frequency distributions. Mean Absolute Error (MAE) of bin-level frequencies quantified the central tendency divergence between real data (R) and generated data (G), calculated as the sum of absolute differences across all 15 bins. Overlap Ratio (OR) measured quantile interval congruence between corresponding bins of R and G distributions. b, Jensen-Shannon Divergence (JSD) comparisons among four datasets: real data (R), baseline generated data (G), China-prompted generated data (GC), and USA-prompted generated data (GA). The inclusion of geographical constraints, such as ’China’, increased GPT-4o’s constraint-triggering rate by 24% (24/100 trials) (e.g., ’I can’t provide real-time data such as the current population’).
@@ 2.4 The potential for improving AI-generated data
A post-hoc calibration procedure based on Optimal Transport (OT) is shown to be the most effective intervention in our comparative study for enhancing data fidelity (Supplementary Experiment 8). Our proposed OT method demonstrates consistent improvements over the raw generative output, reducing the Jensen-Shannon Divergence (JSD) for the GDP distribution from approximately 0.6 to 0.35 and increasing the Overlapping Ratio (OR) for infrastructure by a factor of five (Supplementary Experiment 8a). More critically, the calibration rectifies the systematic parametric biases found in the raw generative output. For instance, it corrects the Zipf’s Law exponent from a biased value of approximately -0.5 to a value that aligns with the theoretical -1.0, and similarly adjusts the scaling exponents for infrastructure and GDP to match their empirical benchmarks (Supplementary Experiment 8b). The method’s effectiveness is further validated by an ablation study, which shows that applying OT to uninformed noise (Noise + OT) yields only limited improvements, confirming the synergistic importance of the GenAI’s sophisticated generative prior. Furthermore, the OT consistently outperforms the traditional statistical synthesizer (KDE), positioning it as a robust method for generating high-fidelity urban data.
2.5 The impact of different prompts and models
Given the observed deviations, this study investigated whether these gaps could be bridged through different forms of intervention. The results indicate that the fidelity of generated data possesses potential for improvement through both user interaction and model selection. Strategic prompting offers a viable pathway to enhance data fidelity. Prompting with specific geographic contexts demonstrates a clear effect on data distribution; when prompted with ”China,” the generated data showed a greater alignment with real data from Chinese cities, increasing the Overlap Ratio by 38.56% and decreasing the Mean Absolute Error by 52.14% (Fig. 5a, 5b). Similarly, at the meso-scale, prompting with parameters like city size or historical period directly influences the resulting decay curves (Fig. 2c). These results demonstrate that the model’s output is sensitive to contextual cues, highlighting a clear potential for iterative refinement.
Furthermore, performance varies noticeably across the different GenAI models evaluated. In the urban scaling law test, GPT-4o demonstrated the closest alignment with empirical results, while Claude 3.5 was most accurate in reproducing Zipf’s law. For distance decay, GPT-4o again showed more nuanced spatial reasoning, whereas models like ChatGLM-4 and DeepSeek-V3 tended to produce less diverse outputs. In the urban vitality assessment, GPT-4o consistently outperformed all other models. These variations suggest that factors intrinsic to the models, such as architecture, training data, and fine-tuning strategies, significantly influence the characteristics and fidelity of the generated urban data.